### v1

# PDSS: A Privacy-Preserving Framework for Step-by-Step Distillation of Large Language Models

Abstract
In the context of real-world applications, leveraging large language models (LLMs) for domain-specific tasks often faces two major challenges: domain-specific knowledge privacy and constrained resources. To address these issues, we propose PDSS, a privacy-preserving framework for step-by-step distillation of LLMs. PDSS works on a server-client architecture, wherein client transmits perturbed prompts to the server’s LLM for rationale generation. The generated rationales are then decoded by the client and used to enrich the training of task-specific small language model(SLM) within a multi-task learning paradigm. PDSS introduces two privacy protection strategies: the Exponential Mechanism Strategy and the Encoder-Decoder Strategy, balancing prompt privacy and rationale usability. Experiments demonstrate the effectiveness of PDSS in various text generation tasks, enabling the training of task-specific SLM with enhanced performance while prioritizing data privacy protection.
PDSS: A Privacy-Preserving Framework for Step-by-Step Distillation of Large Language Models
Tao Fan1, 2, Yan Kang2, Weijing Chen2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1 Hong Kong University of Science and Technology, China 2 WeBank, China Correspondence: tfanac@cse.ust.hk
1 Introduction
Large Language Models(LLMs), boasting billions of parameters and remarkable text generation abilities, have risen as a revolutionary force in artificial intelligence. Prominent models, such as GPT-4 OpenAI (2023), LLaMATouvron et al. (2023), and QwenBai et al. (2023), have garnered the attention of researchers and practitioners alike, demonstrating unparalleled proficiency across numerous tasks. Nevertheless, the sheer size of these models presents significant obstacles for real-world deployment, particularly in environments with limited resources. Meanwhile, as LLMs gain escalating popularity and widespread utilization, privacy concerns have moved to the forefront, especially when it comes to user data and model inference. In contrast, Small Language Models(SLMs) often exhibit superior computational efficiency and faster convergence rates, rendering them perfectly suited for real-time applications or resource-constrained environments. Nonetheless, SLMs also possess certain drawbacks stemming from their performance limitations. The question then arises: How can we effectively combine the predictive prowess of LLMs with the nimbleness of SLMs, all while adhering to privacy requirements?
To address these challenges, we introduce PDSS, a privacy-preserving framework for step-by-step distillation of LLMs. In our envisioned setup, there’s a high-powered server capable of deploying an LLM, paired with a client possessing more limited computational resources running SLM. The challenge lies in maintaining the privacy of client data while leveraging the server’s LLM to aid in training the client’s SLM for text generation tasks, thereby elevating its performance. PDSS aims to bridge this gap, enabling secure and efficient knowledge transfer between LLM and SLM, and ultimately enhancing the capabilities of the SLM without compromising privacy.
As illustrated in Figure 1, within our framework, the process works as follows. Initially, the client transmits perturbed prompts to the server’s LLM, which are protected by the PDSS prompt encoder module, thus ensuring privacy protection. Subsequently, the server’s LLM generates perturbed rationales from these prompts through the Chain of Thought (COT) approach Wei et al. (2022) and relays them back to the client. Upon receiving these perturbed rationales, the client’s rationales decoder module reconstructs them into their original, aligned form corresponding to the raw prompt. Ultimately, the client incorporates these rationales as supplementary and enriching information for training its Task-Specific SLM within a multi-task learning paradigm Wei et al. (2022); Hsieh et al. (2023); Zhang and Yang (2021). These rationales justify the predicted labels and serve as insightful guidance for training smaller and domain-specific models.
Within the PDSS framework, to achieve a balance between preserving the privacy of user prompts and enhancing the usability of rationales, we introduce two privacy protection strategies incorporated into the the prompt encoder module and the rationales decoder module: the Exponential Mechanism Strategy and the Encoder-Decoder Strategy. In the Exponential Mechanism Strategy, we utilize an exponential mechanism to obfuscate the prompts Tong et al. (2023), followed by decoding the perturbed rationales through In-Context Learning (ICL) Dong et al. (2022). In the Encoder-Decoder strategy, we utilize an Encoder-Decoder SLM specifically designed to encode raw prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form. To effectively train this unified Encoder-Decoder SLM, we utilize a multi-task learning paradigm Zhang and Yang (2021), encompassing both the encoding and decoding training processes.
Our contributions are summarized as follows:
  • Privacy-Preserving Framework for LLM Distillation. We propose PDSS, a novel framework that facilitates secure and efficient knowledge transfer from LLM to SLM in resource-constrained environments while adhering to privacy requirements. PDSS addresses the challenges posed by the massive size of LLMs for real-world deployment and the privacy concerns surrounding user data. By utilizing perturbed prompts and rationales, PDSS ensures data privacy while leveraging the predictive prowess of LLMs to enhance the performance of SLMs.
  • Innovative Privacy Protection Strategies. Within PDSS, we introduce two privacy protection strategies: the Exponential Mechanism Strategy and the Encoder-Decoder Strategy. The former utilizes an exponential mechanism to obfuscate user prompts, while the latter employs a specialized Encoder-Decoder SLM to encode and decode perturbed prompts and rationales. These strategies effectively balance user privacy and the usability of rationales, allowing for secure and enhanced training of the client’s SLM without compromising on privacy concerns.
  • Empirical Evaluation and Enhanced Performance of Task-Specific SLM. Through experiments on various text generation tasks, PDSS demonstrates the effectiveness of its framework in training task-specific SLM with enhanced performance. By harnessing the rationales generated by the server-side LLM, PDSS provides valuable task-specific knowledge to the SLM, enabling them to achieve significant improvements with the support of the LLM while prioritizing data privacy protections.
Figure 1: Overview of our proposed PDSS workflow.
Figure 2: Privacy-Preserving Rationals Generation Example.
2 Related Work
2.1 Chain of Thought in Large Language Models
The Chain of Thought(COT) approach has recently garnered significant attention in the realm of LLMs, thanks primarily to its remarkable ability to enhance the reasoning capabilities of these models. This innovative concept was first introduced by Wei et al. (2022). Their research demonstrated that by prompting LLMs to produce a sequence of intermediary reasoning steps(rationales), the models’ performance in handling intricate reasoning tasks could be notably boosted. This groundbreaking study opened the door for further explorations into COT. Since the introduction of COT, several studies have delved into its extensions and variations. For example, Kojima et al. (2022) proposed the use of zero-shot COT, where the model is prompted to generate reasoning steps(rationales) without relying on prior examples. COT has also been applied to various domains, including arithmetic reasoningCobbe et al. (2021), commonsense reasoningKlein and Nabi (2020).
Nonetheless, despite the impressive feats achieved by LLMs, the adoption of LLMs in domain-specific applications with constrained resources poses a significant challengeFan et al. (2023) Kang et al. (2023). Recent studies by Hsieh et al. (2023) Ho et al. (2022) Li et al. (2023), have capitalized on the generated rationales as a form of insightful supervision to train smaller and domain-specific models. However, previous studies have not addressed the domain-specific data privacy issue that arises when LLMs and domain-specific smaller models are deployed across different parties. In our work, we endeavor to address this significant challenge.
2.2 Privacy Preserving LLM Inference
With the escalating popularity and widespread utilization of LLMs, privacy concerns have taken center stage, particularly regarding user data and model inference. Previous research efforts aimed at preserving privacy during LLM inference have predominantly focused on several key techniques, including differential privacy(DP) Dwork (2006), fully homomorphic encryption(FHE) Gentry (2009), and secure multi-party computation(MPC) Yao (1986) protocols.
Numerous studies have delved into the intricacies of LLM inference leveraging DP techniques. Notably, methods like SANTEXT+ Yue et al. (2021), CUSTEXT+ Chen et al. (2022), TextObfuscator Zhou et al. (2023) and InferDPT Tong et al. (2023) have harnessed differential privacy to sequentially replace sensitive words in the text with semantically similar alternatives from a predefined word adjacency list.
FHE and MPC techniques have also garnered attention as viable methods for ensuring privacy during LLM inference. For instance, CipherGPT Hou et al. (2023) proposes a secure matrix multiplication and a novel protocol for securely computing GELU within transformer architecture using FHE and MPC protocols to facilitate secure two-party GPT inference. Likewise, Puma Dong et al. (2023) has adopted FHE and MPC in its transformer architecture for secure third-party LLM inference. While FHE and MPC can be utilized for privacy-preserving text generation tasks, their practical applications remain limited primarily due to significant computational and communication overheads.
The advancements in privacy-preserving techniques, such as differential privacy, FHE, and MPC, offer promising solutions to mitigate privacy risks associated with LLM inference. However, balancing privacy and efficiency remains a challenge that requires further exploration and refinement.
3 The Proposed PDSS Framework
3.1 Overview
In this section, we introduce PDSS, an innovative privacy-preserving framework specifically designed for distilling step-by-step LLMs. The PDSS framework can enhance the performance of SLMs while maintaining privacy, leveraging the capabilities of LLM. We illustrate the PDSS in Figure 1 and describe the associated training algorithm in Algorithm 1. The workflow of PDSS is outlined as follows:
  1. 1.
    In the client, Prompt Encoder Module perturbs these prompts before sending them to the server-side LLM.
  2. 2.
    In the server, the server-side LLM generates perturbed rationales based on these perturbed prompts and sends them back to the client.
  3. 3.
    In the client, Rationales Decoder Module decodes the perturbed rationales.
  4. 4.
    In the client, Task-Specific SLM Training Module employs both the original label data and the filter rationales data for multi-task learning.
3.2 Prompt Encoder Module
In the prompt encoder module, as illustrated in Figure 3, we propose two privacy protection strategies:
  1. 1.
    Exponential Mechanism Encoder Strategy. In the first strategy, we utilize an exponential mechanism McSherry and Talwar (2007)Tong et al. (2023), which satisfies the criteria for the $\epsilon-DP$. This strategy works by replacing each token in the prompt with a semantically similar one sampled from either a predetermined adjacency list or a randomly generated adjacency list, based on exponential mechanism.
    The Definition of Exponential Mechanism Tong et al. (2023). For a given scoring function $u:X\times Y\to R$, a randomized mechanism $M(X,u,Y)$ is $\epsilon-DP$ compliant if it satisfies:
    $P_{r}[y|x]\propto exp(\frac{\epsilon\cdot u(x,y)}{2\bigtriangleup u})$ (1)
    where the sensitivity $\bigtriangleup u$ is defined as:
    $\bigtriangleup u=\max_{x,x^{{}^{\prime}}\in X,y\in Y}|u(x,y)-u(x^{{}^{\prime}},y)|$ (2)
  2. 2.
    Encoder-Decoder Encoder Strategy. The tokens within a prompt differ significantly in terms of their importance and degree of privacy. Applying a uniform privacy budget $\epsilon$ across all tokens may not lead to the most optimal solution. To further optimize the privacy-utility balance, we propose an Encoder-Decoder strategy. This strategy is built upon the first exponential mechanism. In the Encoder-Decoder strategy, we utilize an Encoder-Decoder SLM specifically designed to encode raw prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form. This strategy involves two training process: encoding training process and decoding training process. In this section, we mainly focus on encoding training process, as illustrated in Figure 3.
    Initially, an encoding training process is required for the Encoder-Decoder SLM. Formally, let’s denote a public dataset as $P=\left\{(p_{i},p^{\epsilon}_{i}))\right\}^{N}_{i=1}$, where $p_{i}$ represents raw private prompt, $p^{\epsilon}_{i}$ represents perturbed prompt generated using the first exponential mechanism with a privacy budget of $\epsilon$. In the encoding training process, we train the Encoder-Decoder SLM: $g_{\phi}(p_{i})\to p^{\epsilon}_{i}$. The details of encoding training process is illustrated in Algorithm 1.
    The Encoder objective can be formulated as follows:
    $\mathcal{L}_{\text{Encoder}}(\phi;\mathcal{P})=\mathbb{E}_{(p,p^{\epsilon})\sim\mathcal{P}}\ell_{\text{CE}}(g_{\phi}(p),p^{\epsilon})$ (3)
    where $\ell_{\text{CE}}$ is the cross-entropy loss.
As illustrated in Figure 2, we can observe an exemplary comparison between the original input and its perturbed input in Step 1 and Step 2. This perturbed prompt serves as the new, privacy-enhanced input for further processing.
By incorporating this perturbation mechanism, we ensure that the privacy of the original prompt is preserved. This approach not only satisfies the privacy requirements but also enables effective data utilization for downstream tasks, striking a balance between privacy and utility.
Figure 3: Prompt Encoder Module.
3.3 Generating Perturbed Rationales from LLM
When the server-side LLM receives the perturbed prompt, we leverage the Chain-of-Thought (CoT) prompting technique introduced by Wei et al. (2022) to generate rationales from the LLM using this perturbed prompt. These generated rationales, which are also perturbed, are then transmitted to the client. For instance, as illustrated in Figure 2, given a perturbed prompt in the Step 2, the LLM generates perturbed rationales in the Step 3.
3.4 Rationales Decoder Module
Once the client receives the perturbed rationales from the server-side LLM, it must initiate a "decoder" process within the rationales decoder module to decode the rationales. In rationales decoder module, as illustrated in Figure 4, we also propose two strategies correspond to the two protection strategy of the prompt encoder module:
  1. 1.
    Exponential Mechanism Decoder Strategy. In the first decoding strategy, which corresponds to Exponential Mechanism Encoder strategy. Here, we utilize In-Context Learning(ICL) Dong et al. (2022) Tong et al. (2023) with the SLM to decode the perturbed rationales. we can input a sample $x_{i}=(p,p^{p},r^{p})_{i}$ into the SLM to prompt the generation of rationales, where $p$ represents raw private prompt, $p^{p}$ represents perturbed prompt and $r^{p}$ represents perturbed rationales generated from LLM. $(p^{p},r^{p})_{i}$ can be viewed as an example for SLM in ICL. This allows the SLM to generate rationales $r_{i}$ that are aligned with the original, unperturbed prompt.
  2. 2.
    Encoder-Decoder Decoder Strategy. In the second decoding strategy, which corresponds to Encoder-Decoder Encoder strategy. The rationales decoder module also use the same the Encoder-Decoder SLM with Section 3.2.
    Initially, a decoding training process is required for the Encoder-Decoder SLM. Formally, let’s denote a public dataset as $R=\left\{(x_{i},r_{i}))\right\}^{N}_{i=1}$, where $x_{i}$ represents an input, where $x_{i}=(p,p^{p},r^{p})_{i}$ , $p$ represents raw private prompt, $p^{p}$ represents perturbed prompt generated from Encoder-Decoder SLM, $r^{p}$ represents perturbed rationales generated from LLM. $r_{i}$ represents the raw rationale of raw prompt $p$ generated from LLM. In the decoding training process, we train the Encoder-Decoder SLM: $g_{\phi}(x_{i})\to r_{i}$. The details of decoding training process is illustrated in Algorithm 1.
    The Decoder objective can be formulated as follows:
    $\mathcal{L}_{\text{Decoder}}(\phi;\mathcal{R})=\mathbb{E}_{(x,r)\sim\mathcal{R}}\ell_{\text{CE}}(g_{\phi}(x),r)$ (4)
    where $\mathcal{L}_{\text{Decoder}}$ is the rational decoder loss, and $\ell_{\text{CE}}$ is the cross-entropy loss.
    Subsequently, once the decoding training process of Encoder-Decoder SLM is finished, we can input a sample $x_{i}=(p,p^{p},r^{p})_{i}$ into the SLM, where $r^{p}$ represents perturbed rationales generated from LLM. This allows the SLM to generate rationales $r_{i}$ that are aligned with the original, unperturbed prompt.
    We approach the training of the Encoder-Decoder SLM as a multi-task learning problem encompassing both the encoding and decoding training processes. The multi-task learning objective can be formulated as follows:
    $\mathcal{L}_{1}=\alpha\mathcal{L}_{\text{Encoder}}+(1-\alpha)\mathcal{L}_{\text{Decoder}}$ (5)
    where $\alpha$ is the hyperparameters that control the weight of encoder and decoder loss.
As illustrated in Figure 2, we can observe an exemplary comparison between the perturbed rationales from LLM and its decoded rationales from SLM in Step 3 and Step 4. It’s worth noting that although the SLM has the ability to generate aligned rationales independently, the quality often falls short due to its limited capabilities. By leveraging the perturbed rationales, we effectively transfer the powerful capabilities of the server-side LLM to enhance the Encoder-Decoder SLM, thereby improving the overall quality of the generated rationales.
Figure 4: Rationales Decoder Module.
Algorithm 1 PDSS
1:
2:$T$: total number of rounds;
3:$\mathcal{P}$: encoding training datasets;
$\mathcal{R}$: decoding training datasets;
5:$\mathcal{D}$: task-Spec training datasets;
6:$\eta_{\phi}$: learning rate of Encoder-Decoder SLM;
7:$\eta_{\omega}$: learning rate of Task-Specific SLM.
8:$g_{\phi}$, $f_{\omega}$.
9$\triangleright$ Multi-Task Training for Encoder-Decoder SLM based on Public Datasets $\mathcal{P}$ and $\mathcal{R}$.
10:for each epoch $t\in[T]$ do
11:     $\phi^{t+1}\leftarrow\phi^{t}-\eta_{\phi}\bigtriangledown\mathcal{L}_{1}$.
12:end for
13$\triangleright$ Generated $p^{p}$ using the updated Encoder.
14:$p^{p}=SLM_{Encoder}(p)$.
15$\triangleright$ Generated perturbed rationales from LLM on the server.
16:$r^{p}=LLM(p^{p})$.
17$\triangleright$ Decoded perturbed rationales using the updated Encoder-Decoder SLM.
18:$r=SLM_{Decoder}(r^{p})$.
19$\triangleright$ Multi-Task Training for Task-Specific SLM based on Datasets $\mathcal{D}$.
20:for each epoch $t\in[T]$ do
21:     $\omega^{t+1}\leftarrow\omega^{t}-\eta_{\omega}\bigtriangledown\mathcal{L}_{2}$.
22:end for
3.5 Task-Specific SLM Training Module
In our work, we undertake the training of the client’s Task-Specific SLM tailored for text generation tasks. Initially, we elaborate on the prevalent framework for learning task-specific models. Leveraging this established framework, we enhance it by integrating rationales produced from the rationales decoder module into the training process. Formally, let’s denote a dataset as $D=\left\{(x_{i},(y_{i},r_{i}))\right\}^{N}_{i=1}$, where $x_{i}$ represents an input, $y_{i}$ represents the associated expected output label, and $r_{i}$ is the corresponding desired rationale.
We conceptualize learning with rationales as a multi-task learning problem, as illustrated in Figure 5. Specifically, we train the model $f_{\omega}(x_{i})\to(y_{i},r_{i})$ to accomplish not just the prediction of task labels but also the generation of the corresponding rationales based on textual inputs. This multi-task training ensures that our model not only produces accurate predictions but also provides insightful justifications for its decisions. By doing so, we enhance the transparency and explainability of the model. The multi-task learning objective can be formulated as follows:
$\mathcal{L}_{2}=\beta\mathcal{L}_{\text{Label}}+(1-\beta)\mathcal{L}_{\text{Rationale}}$ (6)
where $\mathcal{L}_{\text{Label}}$ is the label prediction loss:
$\mathcal{L}_{\text{Label}}(\omega;\mathcal{D})=\mathbb{E}_{(x,y)\sim\mathcal{D}}\ell_{\text{CE}}(f_{\omega}(x),y)$ (7)
and $\mathcal{L}_{\text{Rationale}}$ is the rationale generation loss:
$\mathcal{L}_{\text{Rationale}}(\omega;\mathcal{D})=\mathbb{E}_{(x,r)\sim\mathcal{D}}\ell_{\text{CE}}(f_{\omega}(x),r)$ (8)
where $\ell_{\text{CE}}$ is the cross-entropy loss, $f_{\omega}(.)$ is the Task-Specific SLM model, and $\beta$ is the hyperparameters that control the weight of label prediction loss and rationale generation loss.
Figure 5: Task-Specific SLM Training Module.
4 Experiments
4.1 Setup
We have established a scenario to evaluate the performance of the PDSS framework across a range of text generation tasks. This setup involves a client-server architecture, where the client holds two downstream SLMs :an Encoder-Decoder SLM, which specializes in encoder-decoder functionalities and a Task-Specific SLM, tailored for specific tasks. On the server-side, we host a LLM for more general and powerful text generation capabilities. Specifically, we have chosen Qwen-14BBai et al. (2023) as LLM, while both SLMs are Qwen-0.5BBai et al. (2023). Table 1 outlines the detailed configurations of both the LLM and the SLMs.
Setting Server Client Client
Model Type LLM Encoder-Decoder SLM Task-Specific SLM
Model Name Qwen-14B Qwen-0.5B Qwen-0.5B
Parameters(Billion) 14 0.5 0.5
Table 1: LLM and SLMs Setting of PDSS.
Datasets and Evaluation Metrics. We conduct a comprehensive evaluation of PDSS on 4 QA datasets. Specifically, we include CommonsenseQA(CQA) Talmor et al. (2018), OpenBookQA(OBQA) Mihaylov et al. (2018), BoolQ Clark et al. (2019), ARC-EClark et al. (2018). For these datasets, we primarily use Accuracy as the evaluation metric.
Baselines. Since we incorporate two distinct strategies in the prompt encoder module and rationales decoder module, we denote PDSS method with the Exponential Mechanism Strategy as PDSS-EM and PDSS method with the Encoder-Decoder Strategy as PDSS-ED. We conduct a comparative analysis to evaluate the performance of our PDSS framework, which comprises both PDSS-EM and PDSS-ED.
These baselines included:
  • FewShot-LLM, which represents the few-shot capabilities of LLM on the server;
  • FewShot-SLM, which represents the few-shot performance of SLM on the client;
  • Standalone, where the client independently fine-tunes its local model using its own private dataset;
  • DSSHsieh et al. (2023), where the client fine-tunes its local model by distilling step-by-step LLM method without privacy-preserving.
4.2 Overall Performance Evaluation
In this section, we undertake a comprehensive analysis of the task performance of PDSS. We assess both the PDSS-EM and PDSS-ED methods against other baselines on Task-Specific SLM across various privacy budgets, denoted by $\epsilon$.
The results, as presented in Table 2, clearly illustrate that both PDSS-EM and PDSS-ED exhibit significantly better performance when compared to FewShot-SLM and Standalone methods. With an increase in the privacy budget $\epsilon$, both the performance of PDSS-EM and PDSS-ED have risen notably. Furthermore, PDSS-ED demonstrates notably superior performance compared to PDSS-EM under the same privacy budget $\epsilon$ Specifically, under a privacy budget of $\epsilon=3$, PDSS-EM surpasses the Standalone method by 3.4% and 17% in the CQA and OBQA datasets, respectively, while PDSS-ED outperforms it by 5.2% and 22.4%. Similarly, when the privacy budget is increased to $\epsilon=10$, PDSS-EM exceeds the Standalone approach by 6.3% and 21.6% within the CQA and OBQA datasets, respectively, and PDSS-ED beats it by 7.2% and 28.6%. Remarkably, across all datasets evaluated, when the privacy budget is set to $\epsilon=10$, PDSS achieves comparable performance to DSS, highlighting its efficacy and versatility in balancing privacy and utility.
Method CQA OBQA BoolQ ARC-E
FewShot-LLM 80.9 82.8 85.2 80.3
FewShot-SLM 25.7 28.6 59.7 40.7
Standalone 55.7 43.4 78.4 50.3
DSS 59.3 55.1 80.5 57.6
PDSS-EM($\epsilon=1$) 57.7 49.2 80.1 52.3
PDSS-EM($\epsilon=3$) 57.6 50.8 79 52.6
PDSS-EM($\epsilon=5$) 58.8 53.2 80 55.3
PDSS-EM($\epsilon=10$) 59.2 52.8 80.2 56.2
PDSS-ED($\epsilon=1$) 58.2 50.8 80.3 56.4
PDSS-ED($\epsilon=3$) 58.6 53.1 80.2 56.5
PDSS-ED($\epsilon=5$) 58.3 53.4 80.4 56.3
PDSS-ED($\epsilon=10$) 59.7 55.8 80.7 57.9
Table 2: We compare the performance of Task-Specific SLM trained with PDSS-EM and PDSS-ED across different privacy budgets $\epsilon$ against the Task-Specific SLM trained using baseline methods.
4.3 Reducing Training Data Evaluation
In this section, we conduct an in-depth analysis to explore the influence of training data size on model performance. We compare the PDSS method with the Standalone approach, varying the amount of training data used. Table 3 provides a clear illustration of how PDSS(with $\epsilon=3$) consistently outperforms the Standalone method.
Remarkably, PDSS achieves superior performance even with significantly fewer training samples compared to Standalone. More specifically, when trained on merely 75% of the complete CQA, OBQA, and BoolQ datasets, both PDSS-EM and PDSS-ED surpasses the performance of Standalone fine-tuning that has been trained on the entirety of these datasets. Likewise, by using only 50% of the full ARC-E dataset, PDSS-EM exceeds the results achieved by Standalone fine-tuning on the complete dataset. Furthermore, PDSS-ED exhibits significantly better performance than PDSS-EM across various dataset sizes (ranging from 25% to 100%). The results indicate that PDSS is capable of extracting more valuable information from smaller datasets, making it a promising approach in data-scarce environments.
Task Method 25% 50% 75% 100%
CQA PDSS-EM 49 53.5 56.7 57.6
PDSS-ED 54.2 54.6 56.1 58.6
Standalone - - - 55.7
OBQA PDSS-EM 34.8 42.2 45.6 50.8
PDSS-ED 41.4 43.6 50.6 53.1
Standalone - - - 44.2
BoolQ PDSS-EM 63 74 78.7 79
PDSS-ED 72.8 77.6 79.1 80.2
Standalone - - - 78.4
ARC-E PDSS-EM 45.3 52.2 53.1 53.8
PDSS-ED 48 49.7 55.9 56.5
Standalone - - - 50.3
Table 3: We compare the performance of Task-Specific SLM trained with PDSS-EM($\epsilon=3$) and PDSS-ED($\epsilon=3$) against Standalone, across a range of dataset sizes from 25% to 100%. The ’-’ indicates a method does not apply to the corresponding dataset sizes.
4.4 Perturbed Rationales Evaluation
In this section, we focus on analyzing the quality of the perturbed rationales($r^{p}$) generated from the perturbed prompt of LLM based on PDSS-EM and PDSS-ED methods and compare them with the rationales($r$) generated from raw prompt of the LLM. To evaluate the similarity between $r^{p}$ and $r$, we use TokenRatio metric. A higher TokenRatio indicates a greater degree of similarity between the perturbed and original rationales. For more details about TokenRatio, please refer to Appendix C.
As shown in Table 4, with an increase in the privacy budget $\epsilon$ and a corresponding decrease in perturbation, both the TokenRatio of PDSS-EM and PDSS-ED have risen notably. Furthermore, in most of tasks, the TokenRatio of PDSS-ED is higher than that of PDSS-EM in the same level of privacy budget $\epsilon$. The experimental results confirm that the TokenRatio observed in the perturbed rationales produced by both PDSS-EM and PDSS-ED, positively correlate with the privacy budget $\epsilon$. This suggests that as the privacy constraints are relaxed (higher $\epsilon$ values), the perturbed rationales become more similar to the original rationales. This finding is significant as it demonstrates the trade-off between privacy protection and the utility of the generated rationales.
Method CQA OBQA BoolQ ARC-E
PDSS-EM($\epsilon=1$) 19.8 26.2 26.6 24.6
PDSS-EM($\epsilon=3$) 29.2 37.2 35.5 33.9
PDSS-EM($\epsilon=5$) 48.8 59.6 55.2 53.9
PDSS-EM($\epsilon=10$) 69.7 72 74.6 68.2
PDSS-ED($\epsilon=1$) 26.7 33.1 29.7 31
PDSS-ED($\epsilon=3$) 33.1 40.9 40.4 42.9
PDSS-ED($\epsilon=5$) 49.6 61 57.5 63.5
PDSS-ED($\epsilon=10$) 57.2 68.3 68 74.2
Table 4: We conduct a comparative analysis to assess the perturbed rationales produced by PDSS-EM and PDSS-ED methods against the original, unperturbed (raw) rationales that are directly generated from the raw prompt of the LLM.
4.5 Decoded Rationales Evaluation
In this section, we delve into the quality analysis of the decoded rationales produced by the rationales decoder module based on PDSS-EM and PDSS-ED methods. We compare these decoded rationales against those generated directly from raw prompt of the LLM. We utilize the TokenRatio metric to assess their similarities.
As shown in Table 5, in contrast to FewShot-SLM, it becomes apparent that the decoded rationales’ quality based on PDSS-EM and PDSS-ED methods isn’t solely reliant on the locally decoded SLM. The perturbed rationales crafted by the LLM indeed fulfill their intended purpose. When juxtaposed with Table 4, it’s clear that at comparable $\epsilon$ levels, the TokenRatio for the decoded rationales surpass those of the perturbed rationales in the PDSS-EM and PDSS-ED methods. This underscores the effectiveness of the rationales decoder module in the PDSS-EM and PDSS-ED methods. Furthermore, with the increase of the privacy budget $\epsilon$, the TokenRatio for the decoded rationales generated by both PDSS-EM and PDSS-ED have increased significantly. This suggests that as the privacy constraints are relaxed (higher $\epsilon$ values), the decoded rationales become more similar to the original rationales. For more details about comparative analysis of perturbed rationales and decoded rationales, please refer to Appendix D.
Method CQA OBQA BoolQ ARC-E
FewShot-SLM 43.3 43.4 51.9 42.6
PDSS-EM($\epsilon=1$) 38.3 37.1 38.4 41.5
PDSS-EM($\epsilon=3$) 41.9 41.3 41.7 45.6
PDSS-EM($\epsilon=5$) 53.1 54 55 58.3
PDSS-EM($\epsilon=10$) 71.1 63 73.6 70.4
PDSS-ED($\epsilon=1$) 57.2 53.4 45.2 57.5
PDSS-ED($\epsilon=3$) 59 55.1 48 59.4
PDSS-ED($\epsilon=5$) 59.8 59.5 55.7 65.5
PDSS-ED($\epsilon=10$) 62 62.3 63.4 70.1
Table 5: We conduct a comparative analysis to assess the decoded rationales produced by PDSS-EM and PDSS-ED methods against the original, unperturbed (raw) rationales that are directly generated from the raw prompt of the LLM.
5 Conclusions
We introduced PDSS, a privacy-preserving framework for LLM distillation, addressing domain-specific knowledge privacy and resource constraints. PDSS employs a server-client architecture with prompt encoding, rationale generating, rationale decoding, and task-specific SLM training, bridging the gap between LLM and SLM while maintaining data privacy. Experiments on various text generation tasks demonstrate PDSS’s ability to enhance SLM performance with LLM support while prioritizing data privacy.
Limitations
Our current study faces limitations due to computational and storage constraints, which hinder our ability to experiment with larger model sizes. Additionally, our evaluation of PDSS has been restricted to the Qwen model architecture, leaving the possibility that PDSS may need to be further explored in other model architectures. We intend to tackle these issues in future research endeavors.
Appendix A Rationales Generation through COT
We utilize the rationales data generated by server-side LLM through chain-of-thought (CoT)Wei et al. (2022)Hsieh et al. (2023) technique to enhance the performance of the client’s task-specific SLM. These rationales justify the predicted labels and serve as insightful guidance for training smaller and domain-specific models. Consider the following example: when asked “Question:A beaver is know for building prowess, their supplies come from where? Answer Choices: (a) british columbia (b) body of water (c) wooded area (d) pay debts (e) zoo”. Utilizing the chain-of-thought (CoT) technique, the LLM can generate intermediate rationales like, "The answer must be the place where beavers get their supplies. Of the above choices, only wooded areas have the supplies that beavers need.” Such rationales bridge the gap between the input and the final answer, often encapsulating valuable task-related knowledge. This knowledge would traditionally require extensive data for smaller and task-specific models to acquire. Therefore, we harness these rationales as enriched training material for small language models, employing a multi-task training paradigm that encompasses both label prediction task and rationale prediction task.
Appendix B More on Experimental Details
B.1 Hyperparameter Settings
SLM Parameters. During the training process for both the Encoder-Decoder SLM and the Task-Specific SLM, we specifically configured the parameters. We set the batch size to 32 and employed the AdamW optimizer. The maximum number of training steps ranged from 400 to 1500. Additionally, we assigned the values of 0.5 to both $\alpha$ and $\beta$. Furthermore, the learning rates for $\eta_{\phi}$ and $\eta_{\omega}$ were established at 5e-5.
B.2 Data Splitting
For the datasets CQA/OBQA/BoolQ//ARC-E/, all splits (training, validation, and test) were downloaded from HuggingFace Lhoest et al. (2021). During the training of the Encoder-Decoder SLM, we randomly divided the training data into two equal parts. One part was designated as the public dataset, while the other part was allocated as the private dataset for the client.
B.3 Dataset Licenses
For the datasets CQA/OBQA/BoolQ//ARC-E/ were downloaded from HuggingFaceLhoest et al. (2021) and under Apache License, Version 2.0.
B.4 Machine Configuration
The experiments were conducted on machines equipped with 4 Nvidia V100 32G.
Appendix C The Definition of TokenRatio Metric
TokenRatio($r^{{}^{\prime}},r$). This metric calculates the unique words($u$) in $r^{{}^{\prime}}$ and counts how many of these words are also present in $r$, denoted as $i$. The TokenRatio is then calculated as $i$ divided by the total number of unique words in $r^{{}^{\prime}}$ ($|u|$).
Figure 6: Comparative Analysis of Perturbed Rationales and Decoded Rationales.
Appendix D Comparative Analysis of Perturbed Rationales and Decoded Rationales
As shown in Figure 6, we conduct a comparison of the quality between the perturbed rationales and the decoded rationales, employing both the PDSS-EM and PDSS-ED methods across various privacy budgets denoted by $\epsilon$. For clarity, we designate the perturbed rationales generated using the PDSS-EM and PDSS-ED methods as P-PDSS-EM and P-PDSS-ED, respectively. Similarly, the decoded rationales derived from these methods are denoted as D-PDSS-EM and D-PDSS-ED. It’s clear that at comparable $\epsilon$ levels, the TokenRatio for decoded rationales consistently surpasses that of perturbed rationales in most tasks, when utilizing the PDSS-EM and PDSS-ED methods. This finding underscores the remarkable effectiveness of the rationales decoder module within both the PDSS-EM and PDSS-ED frameworks.

### v2

@@
-PDSS: A Privacy-Preserving Framework for Step-by-Step Distillation of Large Language Models
+FedCoT: Federated Chain-of-Thought Distillation for Large Language Models
 Abstract
-In the context of real-world applications, leveraging large language models (LLMs) for domain-specific tasks often faces two major challenges: domain-specific knowledge privacy and constrained resources.
-To address these issues, we propose PDSS, a privacy-preserving framework for step-by-step distillation of LLMs.
-PDSS works on a server-client architecture, wherein client transmits perturbed prompts to the server’s LLM for rationale generation.
-The generated rationales are then decoded by the client and used to enrich the training of task-specific small language model(SLM) within a multi-task learning paradigm.
-PDSS introduces two privacy protection strategies: the Exponential Mechanism Strategy and the Encoder-Decoder Strategy, balancing prompt privacy and rationale usability.
-Experiments demonstrate the effectiveness of PDSS in various text generation tasks, enabling the training of task-specific SLM with enhanced performance while prioritizing data privacy protection.
-PDSS: A Privacy-Preserving Framework for Step-by-Step Distillation of Large Language Models
-Tao Fan1, 2, Yan Kang2, Weijing Chen2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1 Hong Kong University of Science and Technology, China 2 WeBank, China Correspondence: tfanac@cse.ust.hk
+Large Language Models (LLMs) have emerged as a transformative force in artificial intelligence, demonstrating exceptional proficiency across various tasks.
+However, their deployment in resource-constrained environments and concerns over user data privacy pose significant challenges.
+In contrast, Small Language Models (SLMs) offer computational efficiency but often lag in performance.
+To address these issues, we propose FedCoT, a federated framework designed for the Chain-of-Thought (CoT) distillation of knowledge from LLMs to SLMs, while ensuring the preservation of clients’ data privacy.
+FedCoT ensures secure and efficient knowledge transfer from an LLM on a high-powered server to an SLM on a resource-constrained client, while adhering to privacy requirements.
+Leveraging perturbed prompts and rationales generated through the CoT approach, the framework enhances the performance of the client’s SLM without compromising user data privacy within a multi-task learning framework.
+We propose two privacy protection strategies: the Exponential Mechanism Strategy and the Adaptive Exponential Mechanism Strategy, which balance user prompt privacy and the usability of rationales.
+Empirical evaluation on various text generation tasks demonstrates the effectiveness of FedCoT in training task-specific SLMs with enhanced performance while prioritizing data privacy protection.
+Our code has been contributed to the FATE open-source project and is now publicly accessible at https://github.com/FederatedAI/FATE-LLM/tree/main/python/fate_llm/algo/fedcot
+FedCoT: Federated Chain-of-Thought Distillation for Large Language Models
+Tao Fan1, 2, Weijing Chen2, Yan Kang2, Guoqiang Ma2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Qiang Yang3 1 Hong Kong University of Science and Technology, Hong Kong, China 2 WeBank, China 3Hong Kong Polytechnic University, Hong Kong, China Correspondence: tfanac@cse.ust.hk, qyang@cse.ust.hk
 1 Introduction
-Large Language Models(LLMs), boasting billions of parameters and remarkable text generation abilities, have risen as a revolutionary force in artificial intelligence.
-Prominent models, such as GPT-4 OpenAI (2023), LLaMATouvron et al.
-(2023), and QwenBai et al.
-(2023), have garnered the attention of researchers and practitioners alike, demonstrating unparalleled proficiency across numerous tasks.
-Nevertheless, the sheer size of these models presents significant obstacles for real-world deployment, particularly in environments with limited resources.
-Meanwhile, as LLMs gain escalating popularity and widespread utilization, privacy concerns have moved to the forefront, especially when it comes to user data and model inference.
+Large Language Models (LLMs) have risen as a revolutionary force in artificial intelligence.
+Prominent LLMs, such as GPT-4 Gpt-4, LLaMA touvron2023llama, and Qwen bai2023qwen, have garnered the attention of researchers and practitioners alike, demonstrating unparalleled proficiency across numerous tasks.
+Nevertheless, the sheer size of these models presents significant obstacles for real-world deployment, particularly in environments with limited resources fan2025ten; fan2025fedmkt; fan2023fate-llm; kang2023grounding.
+Meanwhile, as LLMs gain escalating popularity and widespread utilization, privacy concerns have moved to the forefront, especially when it comes to user data and LLMs inference.
 In contrast, Small Language Models (SLMs) often exhibit superior computational efficiency and faster convergence rates, rendering them perfectly suited for real-time applications or resource-constrained environments.
@@
 The question then arises: How can we effectively combine the predictive prowess of LLMs with the nimbleness of SLMs, all while adhering to privacy requirements?
-To address these challenges, we introduce PDSS, a privacy-preserving framework for step-by-step distillation of LLMs.
-In our envisioned setup, there’s a high-powered server capable of deploying an LLM, paired with a client possessing more limited computational resources running SLM.
+To address these challenges, we propose FedCoT, a federated framework designed for the Chain-of-Thought (CoT) wei2022chain distillation of knowledge from LLMs to SLMs, while ensuring the preservation of clients’ data privacy.
+FedCoT ensures secure and efficient knowledge transfer from an LLM on a high-powered server to an SLM on a resource-constrained client.
 The challenge lies in maintaining the privacy of client data while leveraging the server’s LLM to aid in training the client’s SLM for text generation tasks, thereby elevating its performance.
-PDSS aims to bridge this gap, enabling secure and efficient knowledge transfer between LLM and SLM, and ultimately enhancing the capabilities of the SLM without compromising privacy.
-As illustrated in Figure 1, within our framework, the process works as follows.
-Initially, the client transmits perturbed prompts to the server’s LLM, which are protected by the PDSS prompt encoder module, thus ensuring privacy protection.
-Subsequently, the server’s LLM generates perturbed rationales from these prompts through the Chain of Thought (COT) approach Wei et al.
-(2022) and relays them back to the client.
-Upon receiving these perturbed rationales, the client’s rationales decoder module reconstructs them into their original, aligned form corresponding to the raw prompt.
-Ultimately, the client incorporates these rationales as supplementary and enriching information for training its Task-Specific SLM within a multi-task learning paradigm Wei et al.
-(2022); Hsieh et al.
-(2023); Zhang and Yang (2021).
+FedCoT aims to bridge this gap, enabling secure and efficient knowledge transfer between LLM and SLM, and ultimately enhancing the capabilities of the SLM without compromising privacy.
+As illustrated in Figure 1(a), within our framework, the process works as follows.
+Initially, the client transmits perturbed prompts to the server’s LLM.
+These prompts are protected by the FedCoT prompt encoder, which employs Differential Privacy (DP) principles dwork2006differential; mcsherry2007mechanism, ensuring privacy protection.
+Subsequently, the server’s LLM generates perturbed rationales from these prompts through the CoT approach and relays them back to the client.
+Upon receiving these perturbed rationales, the client’s rationales decoder reconstructs them into their original, aligned form corresponding to the raw prompt.
+Ultimately, the client utilizes CoT knowledge distillation hsieh2023distilling; li2023symbolic to train its Task-Specific SLM.
+This process leverages both label data and rationales within a multi-task learning paradigm wei2022chain; hsieh2023distilling; zhang2021survey.
 These rationales justify the predicted labels and serve as insightful guidance for training smaller and domain-specific models.
-Within the PDSS framework, to achieve a balance between preserving the privacy of user prompts and enhancing the usability of rationales, we introduce two privacy protection strategies incorporated into the the prompt encoder module and the rationales decoder module: the Exponential Mechanism Strategy and the Encoder-Decoder Strategy.
-In the Exponential Mechanism Strategy, we utilize an exponential mechanism to obfuscate the prompts Tong et al.
-(2023), followed by decoding the perturbed rationales through In-Context Learning (ICL) Dong et al.
-(2022).
-In the Encoder-Decoder strategy, we utilize an Encoder-Decoder SLM specifically designed to encode raw prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form.
-To effectively train this unified Encoder-Decoder SLM, we utilize a multi-task learning paradigm Zhang and Yang (2021), encompassing both the encoding and decoding training processes.
+Previous endeavors to incorporate DP into language models, specifically through DP-SGD song2013stochastic, have primarily centered on navigating the delicate balance between utility and privacy.
+This is achieved by introducing calibrated noise into gradients or text representations during the model training process.
+Nonetheless, these methods inherently rely on a trusted server to gather data from data owners for model training chen-etal-2023-customized, significantly limiting their applicability in scenarios where such trusted servers are not available, as is the case in our research context.
+Within the FedCoT framework, to achieve a balance between preserving the privacy of user prompts and enhancing the usability of rationales, we introduce two privacy protection strategies: the Exponential Mechanism Strategy and the Adaptive Exponential Mechanism Strategy.
+In the Exponential Mechanism Strategy, we utilize an exponential mechanism to obfuscate the prompts mcsherry2007mechanism; yue2021differential; chen-etal-2023-customized, followed by decoding the perturbed rationales through In-Context Learning (ICL) dong2024survey; tong2025inferdpt.
+In the Adaptive Exponential Mechanism Strategy, we utilize an Encoder-Decoder SLM specifically designed to encode original prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form.
+To effectively train this unified Encoder-Decoder SLM, we utilize a multi-task learning paradigm zhang2021survey, encompassing both the encoding and decoding training processes.
 Our contributions are summarized as follows:
-  • Privacy-Preserving Framework for LLM Distillation.
-We propose PDSS, a novel framework that facilitates secure and efficient knowledge transfer from LLM to SLM in resource-constrained environments while adhering to privacy requirements.
-PDSS addresses the challenges posed by the massive size of LLMs for real-world deployment and the privacy concerns surrounding user data.
-By utilizing perturbed prompts and rationales, PDSS ensures data privacy while leveraging the predictive prowess of LLMs to enhance the performance of SLMs.
-  • Innovative Privacy Protection Strategies.
-Within PDSS, we introduce two privacy protection strategies: the Exponential Mechanism Strategy and the Encoder-Decoder Strategy.
-The former utilizes an exponential mechanism to obfuscate user prompts, while the latter employs a specialized Encoder-Decoder SLM to encode and decode perturbed prompts and rationales.
-These strategies effectively balance user privacy and the usability of rationales, allowing for secure and enhanced training of the client’s SLM without compromising on privacy concerns.
+  • Federated Framework for CoT Distillation in LLMs.
+We propose FedCoT, a novel federated framework that facilitates secure and efficient knowledge transfer from LLM to SLM in resource-constrained environments.
+FedCoT leverages CoT knowledge distillation to enhance Task-Specific SLM within the client.
+This process leverages rationales produced by the LLM on the server, thereby enriching the client-side SLMs with valuable task-related knowledge.
+  • Privacy as a Priority.
+FedCoT leverages an Adaptive Exponential Mechanism Strategy tailored for encoding prompt to ensure their obfuscation and decoding perturbed rationales.
+The strategies effectively balance user prompt privacy and the usability of rationales.
   • Empirical Evaluation and Enhanced Performance of Task-Specific SLM.
-Through experiments on various text generation tasks, PDSS demonstrates the effectiveness of its framework in training task-specific SLM with enhanced performance.
-By harnessing the rationales generated by the server-side LLM, PDSS provides valuable task-specific knowledge to the SLM, enabling them to achieve significant improvements with the support of the LLM while prioritizing data privacy protections.
-Figure 1: Overview of our proposed PDSS workflow.
-Figure 2: Privacy-Preserving Rationals Generation Example.
+Through experiments on various text generation tasks, FedCoT demonstrates the effectiveness of its framework in training task-specific SLM with enhanced performance.
+By harnessing the rationales generated by the server-side LLM, FedCoT provides valuable task-specific knowledge to the SLM.
+(a) Overview of our proposed FedCoT framework.
+(b) Privacy-Preserving Rationals Generation.
+Figure 1: The overview of our proposed FedCoT.
+The FedCoT comprises four key components: (1) The Prompt Encoder, which perturbs user prompts to ensure privacy; (2) The LLM, generating perturbed rationales based on the perturbed prompts; (3) The Perturbed Rationales Decoder, which decodes the perturbed rationales back into a usable form; (4) The Task-Specific SLM Enhancing via CoT Knowledge Distillation, utilizing both original labeled data and filtered rationales data for multi-task learning.
 2 Related Work
+2.1 Differential Privacy
+In this section, We briefly revisit two important definitions of differential privacy: $\epsilon$-Differential Privacy and Exponential Mechanism (EM).
+$\epsilon$-Differential Privacy.
+Differential privacy (DP)  dwork2006differential is a rigorous mathematical framework that provides strong privacy guarantees for data analysis.
+It ensures that the output of an algorithm remains statistically indistinguishable whether a particular individual’s data is included or excluded from the dataset.
+Formally, a randomized mechanism $M$ provides $\epsilon$-differential privacy if for all neighboring datasets $D$ and $D^{\prime}$ (differing in at most one record) and for all sets $S$ of possible outputs:
+$\Pr[M(D)\in S]\leq e^{\epsilon}\Pr[M(D^{\prime})\in S]$ (1)
+where $\epsilon$ is the privacy budget that controls the level of indistinguishability.
+Exponential Mechanism.
+The Exponential Mechanism  mcsherry2007mechanism allows for the selection of an outcome from a set of possible outcomes with probabilities proportional to the exponential of their utility scores.
+Formally, given a utility function $u:D\times R\rightarrow\mathbb{R}$ that maps each dataset $D$ and possible outcome $r$ to a real-valued score, the Exponential Mechanism $M(D,u,R)$ satisfies $\epsilon$-differential privacy if it selects and outputs an $r\in R$ with probability:
+$\Pr[M(D)=r]\propto\exp\left(\frac{\epsilon u(D,r)}{2\Delta u}\right)$ (2)
+where $\Delta u$ is the sensitivity of the utility function (in our work, we use cosine similarity as the utility function), defined as the maximum change in utility score when a single record is added or removed from the dataset:
+$\Delta u=\max_{D,D^{\prime},r}|u(D,r)-u(D^{\prime},r)|$ (3)
 2.2 Chain of Thought in Large Language Models
 The Chain of Thought (CoT) approach has recently garnered significant attention in the realm of LLMs, thanks primarily to its remarkable ability to enhance the reasoning capabilities of these models.
-This innovative concept was first introduced by Wei et al.
-(2022).
+This innovative concept was first introduced by  wei2022chain.
 Their research demonstrated that by prompting LLMs to produce a sequence of intermediary reasoning steps (rationales), the models’ performance in handling intricate reasoning tasks could be notably boosted.
-This groundbreaking study opened the door for further explorations into COT.
 Since the introduction of CoT, several studies have delved into its extensions and variations.
-For example, Kojima et al.
-(2022) proposed the use of zero-shot COT, where the model is prompted to generate reasoning steps(rationales) without relying on prior examples.
-COT has also been applied to various domains, including arithmetic reasoningCobbe et al.
-(2021), commonsense reasoningKlein and Nabi (2020).
-Nonetheless, despite the impressive feats achieved by LLMs, the adoption of LLMs in domain-specific applications with constrained resources poses a significant challengeFan et al.
-(2023) Kang et al.
-(2023).
-Recent studies by Hsieh et al.
-(2023) Ho et al.
-(2022) Li et al.
-(2023), have capitalized on the generated rationales as a form of insightful supervision to train smaller and domain-specific models.
+For example,  kojima2022large proposed the use of zero-shot CoT, where the model is prompted to generate rationales without relying on prior examples.
+CoT has also been applied to various domains, including arithmetic reasoning cobbe2021training, commonsense reasoning klein2020contrastive.
+Recent studies by hsieh2023distilling; ho2023large; li2023symbolic, have capitalized on the generated rationales as a form of insightful supervision to train smaller and domain-specific models.
 However, previous studies have not addressed the domain-specific data privacy issue that arises when LLMs and domain-specific smaller models are deployed across different parties.
 In our work, we endeavor to address this significant challenge.
-2.2 Privacy Preserving LLM Inference
-With the escalating popularity and widespread utilization of LLMs, privacy concerns have taken center stage, particularly regarding user data and model inference.
-Previous research efforts aimed at preserving privacy during LLM inference have predominantly focused on several key techniques, including differential privacy(DP) Dwork (2006), fully homomorphic encryption(FHE) Gentry (2009), and secure multi-party computation(MPC) Yao (1986) protocols.
-Numerous studies have delved into the intricacies of LLM inference leveraging DP techniques.
-Notably, methods like SANTEXT+ Yue et al.
-(2021), CUSTEXT+ Chen et al.
-(2022), TextObfuscator Zhou et al.
-(2023) and InferDPT Tong et al.
-(2023) have harnessed differential privacy to sequentially replace sensitive words in the text with semantically similar alternatives from a predefined word adjacency list.
-FHE and MPC techniques have also garnered attention as viable methods for ensuring privacy during LLM inference.
-For instance, CipherGPT Hou et al.
-(2023) proposes a secure matrix multiplication and a novel protocol for securely computing GELU within transformer architecture using FHE and MPC protocols to facilitate secure two-party GPT inference.
-Likewise, Puma Dong et al.
-(2023) has adopted FHE and MPC in its transformer architecture for secure third-party LLM inference.
-While FHE and MPC can be utilized for privacy-preserving text generation tasks, their practical applications remain limited primarily due to significant computational and communication overheads.
-The advancements in privacy-preserving techniques, such as differential privacy, FHE, and MPC, offer promising solutions to mitigate privacy risks associated with LLM inference.
-However, balancing privacy and efficiency remains a challenge that requires further exploration and refinement.
-3 The Proposed PDSS Framework
-3.1 Overview
-In this section, we introduce PDSS, an innovative privacy-preserving framework specifically designed for distilling step-by-step LLMs.
-The PDSS framework can enhance the performance of SLMs while maintaining privacy, leveraging the capabilities of LLM.
-We illustrate the PDSS in Figure 1 and describe the associated training algorithm in Algorithm 1.
-The workflow of PDSS is outlined as follows:
-  1. 1.
-    In the client, Prompt Encoder Module perturbs these prompts before sending them to the server-side LLM.
-  2. 2.
-    In the server, the server-side LLM generates perturbed rationales based on these perturbed prompts and sends them back to the client.
-  3. 3.
-    In the client, Rationales Decoder Module decodes the perturbed rationales.
-  4. 4.
-    In the client, Task-Specific SLM Training Module employs both the original label data and the filter rationales data for multi-task learning.
-3.2 Prompt Encoder Module
-In the prompt encoder module, as illustrated in Figure 3, we propose two privacy protection strategies:
+3 The Proposed FedCoT Framework
+In this section, we introduce FedCoT, a federated framework designed for the CoT distillation of knowledge from LLMs hosted on a high-powered server to SLMs deployed on a resource-constrained client.
+The FedCoT framework can enhance the performance of SLMs while maintaining client data’s privacy, leveraging the capabilities of LLM.
+We assume the server to be semi-honest, implying that it may attempt to recover the private data of the client from the information it observes.
+We illustrate the FedCoT in Figure 1(a), outline its training algorithm in Algorithm 1, and detail its resource requirements in Appendix A.
+Algorithm 1 FedCoT
+Input:
+$T$: total number of rounds;
+$\mathcal{P}$: encoding training datasets;
+$\mathcal{R}$: decoding training datasets;
+$\mathcal{D}$: task-specific training datasets;
+$\eta_{\phi}$: learning rate of Encoder-Decoder SLM;
+$\eta_{\omega}$: learning rate of Task-Specific SLM.
+Output: $g_{\phi}$, $f_{\omega}$.
+[…]
+2:for each epoch $t\in[T]$ do
+3:  $\phi^{t+1}\leftarrow\phi^{t}-\eta_{\phi}\nabla\mathcal{L}_{1}$.
+4:end for
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+12:for each epoch $t\in[T]$ do
+13:  $\omega^{t+1}\leftarrow\omega^{t}-\eta_{\omega}\nabla\mathcal{L}_{2}$.
+14:end for
+3.1 Privacy Preserving Prompt Encoder
+Before the client transmits its raw prompts to the server-side LLM, we need the privacy protection strategy to protect the raw prompts. In this section, we propose two privacy protection strategies:
   1. 1.
     Exponential Mechanism Encoder Strategy.
-In the first strategy, we utilize an exponential mechanism McSherry and Talwar (2007)Tong et al.
-(2023), which satisfies the criteria for the $\epsilon-DP$.
-This strategy works by replacing each token in the prompt with a semantically similar one sampled from either a predetermined adjacency list or a randomly generated adjacency list, based on exponential mechanism.
-    The Definition of Exponential Mechanism Tong et al. (2023). For a given scoring function $u:X\times Y\to R$, a randomized mechanism $M(X,u,Y)$ is $\epsilon-DP$ compliant if it satisfies:
-    $P_{r}[y|x]\propto exp(\frac{\epsilon\cdot u(x,y)}{2\bigtriangleup u})$ (1)
-    where the sensitivity $\bigtriangleup u$ is defined as:
-    $\bigtriangleup u=\max_{x,x^{{}^{\prime}}\in X,y\in Y}|u(x,y)-u(x^{{}^{\prime}},y)|$ (2)
+In the first strategy, we utilize an exponential mechanism mcsherry2007mechanism, which satisfies the criteria for the $\epsilon$-DP.
+For detailed information about the exponential mechanism, please refer to Section 2.1.
+    Let us consider an Exponential Mechanism $M(\cdot)$.
+Given a input prompt $p=\{x_{i}\}_{i=1}^{S}$ comprising $S$ tokens, a set $X$ encompassing all possible input tokens, and a set $Y$ of all potential output tokens, the mechanism $M(\cdot)$ is applied to each input token $x_{i}\in p$.
+If $x_{i}$ belongs to $X$, it is replaced with an output token $y_{i}$ from $Y$.
+Through this process, we obtain a perturbed prompt $p^{p}=\{y_{i}\}_{i=1}^{S}$.
   2. 2.
-    Encoder-Decoder Encoder Strategy.
+    Adaptive Exponential Mechanism Encoder Strategy.
 The tokens within a prompt differ significantly in terms of their importance and degree of privacy.
 Applying a uniform privacy budget $\epsilon$ across all tokens may not lead to the most optimal solution.
-To further optimize the privacy-utility balance, we propose an Encoder-Decoder strategy.
+To further optimize the privacy-utility balance, we propose an Adaptive Exponential Mechanism Encoder strategy.
 This strategy is built upon the first exponential mechanism.
-In the Encoder-Decoder strategy, we utilize an Encoder-Decoder SLM specifically designed to encode raw prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form.
+In the Adaptive Exponential Mechanism Encoder strategy, we utilize an Encoder-Decoder SLM specifically designed to encode raw prompts into perturbed prompts and subsequently decode perturbed rationales back into their original form.
 This strategy involves two training process: encoding training process and decoding training process.
-In this section, we mainly focus on encoding training process, as illustrated in Figure 3.
+In this section, we mainly focus on encoding training process.
     Initially, an encoding training process is required for the Encoder-Decoder SLM.
@@
 The details of encoding training process is illustrated in Algorithm 1.
-    The Encoder objective can be formulated as follows:
+    The Prompt Encoder objective can be formulated as follows:
     $\mathcal{L}_{\text{Encoder}}(\phi;\mathcal{P})=\mathbb{E}_{(p,p^{\epsilon})\sim\mathcal{P}}\ell_{\text{CE}}(g_{\phi}(p),p^{\epsilon})$ (4)
     where $\ell_{\text{CE}}$ is the cross-entropy loss.
-As illustrated in Figure 2, we can observe an exemplary comparison between the original input and its perturbed input in Step 1 and Step 2.
+As illustrated in Figure1(b), we can observe an exemplary comparison between the original prompt and its perturbed prompt in Step 1 and Step 2.
 This perturbed prompt serves as the new, privacy-enhanced input for further processing.
-By incorporating this perturbation mechanism, we ensure that the privacy of the original prompt is preserved.
-This approach not only satisfies the privacy requirements but also enables effective data utilization for downstream tasks, striking a balance between privacy and utility.
-Figure 3: Prompt Encoder Module.
 3.2 Generating Perturbed Rationales from LLM
-When the server-side LLM receives the perturbed prompt, we leverage the Chain-of-Thought (CoT) prompting technique introduced by Wei et al.
-(2022) to generate rationales from the LLM using this perturbed prompt.
+When the server-side LLM receives the perturbed prompt, we leverage the Chain-of-Thought (CoT) prompting technique introduced by wei2022chain to generate rationales from the LLM using this perturbed prompt.
 These generated rationales, which are also perturbed, are then transmitted to the client.
-For instance, as illustrated in Figure 2, given a perturbed prompt in the Step 2, the LLM generates perturbed rationales in the Step 3.
-3.4 Rationales Decoder Module
-Once the client receives the perturbed rationales from the server-side LLM, it must initiate a "decoder" process within the rationales decoder module to decode the rationales.
-In rationales decoder module, as illustrated in Figure 4, we also propose two strategies correspond to the two protection strategy of the prompt encoder module:
+For instance, as illustrated in Figure 1(b), given a perturbed prompt in the Step 2, the LLM generates perturbed rationales in the Step 3.
+3.3 Perturbed Rationales Decoder
+Once the client receives the perturbed rationales from the server-side LLM, it must initiate a "decoder" process to decode the rationales.
+In this section, we also propose two strategies correspond to the two protection strategy of the prompt encoder module:
   1. 1.
@@
 In the first decoding strategy, which corresponds to Exponential Mechanism Encoder strategy.
-Here, we utilize In-Context Learning(ICL) Dong et al.
-(2022) Tong et al.
-(2023) with the SLM to decode the perturbed rationales.
-we can input a sample $x_{i}=(p,p^{p},r^{p})_{i}$ into the SLM to prompt the generation of rationales, where $p$ represents raw private prompt, $p^{p}$ represents perturbed prompt and $r^{p}$ represents perturbed rationales generated from LLM.
-$(p^{p},r^{p})_{i}$ can be viewed as an example for SLM in ICL.
-This allows the SLM to generate rationales $r_{i}$ that are aligned with the original, unperturbed prompt.
+Here, we utilize In-Context Learning (ICL)  dong2024survey; tong2025inferdpt with the Encoder-Decoder SLM to decode the perturbed rationales.
+we can input a sample $x_{i}=(p,p^{p},r^{p})_{i}$ into the Encoder-Decoder SLM to prompt the generation of rationales, where $p$ represents raw private prompt, $p^{p}$ represents perturbed prompt and $r^{p}$ represents perturbed rationales generated from LLM.
+$(p^{p},r^{p})_{i}$ can be viewed as an example for Encoder-Decoder SLM in ICL.
+This allows the Encoder-Decoder SLM to generate rationales $r_{i}$ that are aligned with the original, unperturbed prompt.
   2. 2.
-    Encoder-Decoder Decoder Strategy.
-In the second decoding strategy, which corresponds to Encoder-Decoder Encoder strategy.
+    Adaptive Exponential Mechanism Decoder Strategy.
+In the second decoding strategy, which corresponds to Adaptive Exponential Mechanism Encoder strategy.
 The rationales decoder module also use the same the Encoder-Decoder SLM with Section 3.1.
@@
 The details of decoding training process is illustrated in Algorithm 1.
-    The Decoder objective can be formulated as follows:
+    The Rationales Decoder objective can be formulated as follows:
     $\mathcal{L}_{\text{Decoder}}(\phi;\mathcal{R})=\mathbb{E}_{(x,r)\sim\mathcal{R}}\ell_{\text{CE}}(g_{\phi}(x),r)$ (5)
-    where $\mathcal{L}_{\text{Decoder}}$ is the rational decoder loss, and $\ell_{\text{CE}}$ is the cross-entropy loss.
     Subsequently, once the decoding training process of Encoder-Decoder SLM is finished, we can input a sample $x_{i}=(p,p^{p},r^{p})_{i}$ into the SLM, where $r^{p}$ represents perturbed rationales generated from LLM.
@@
     We approach the training of the Encoder-Decoder SLM as a multi-task learning problem encompassing both the encoding and decoding training processes.
-The multi-task learning objective can be formulated as follows:
-    $\mathcal{L}_{1}=\alpha\mathcal{L}_{\text{Encoder}}+(1-\alpha)\mathcal{L}_{\text{Decoder}}$ (5)
-    where $\alpha$ is the hyperparameters that control the weight of encoder and decoder loss.
-As illustrated in Figure 2, we can observe an exemplary comparison between the perturbed rationales from LLM and its decoded rationales from SLM in Step 3 and Step 4.
+    The multi-task learning objective for the Encoder-Decoder SLM can be formulated as follows:
+    $\mathcal{L}_{1}=\mathcal{L}_{\text{Encoder}}+\mathcal{L}_{\text{Decoder}}$ (6)
+As illustrated in Figure1(b), we can observe an exemplary comparison between the perturbed rationales from LLM and its decoded rationales from SLM in Step 3 and Step 4.
 It’s worth noting that although the SLM has the ability to generate aligned rationales independently, the quality often falls short due to its limited capabilities.
 By leveraging the perturbed rationales, we effectively transfer the powerful capabilities of the server-side LLM to enhance the Encoder-Decoder SLM, thereby improving the overall quality of the generated rationales.
-Figure 4: Rationales Decoder Module.
-Algorithm 1 PDSS
-1:
-2:$T$: total number of rounds;
-3:$\mathcal{P}$: encoding training datasets;
-$\mathcal{R}$: decoding training datasets;
-5:$\mathcal{D}$: task-Spec training datasets;
-6:$\eta_{\phi}$: learning rate of Encoder-Decoder SLM;
-7:$\eta_{\omega}$: learning rate of Task-Specific SLM.
-8:$g_{\phi}$, $f_{\omega}$.
-9$\triangleright$ Multi-Task Training for Encoder-Decoder SLM based on Public Datasets $\mathcal{P}$ and $\mathcal{R}$.
-10:for each epoch $t\in[T]$ do
-11:     $\phi^{t+1}\leftarrow\phi^{t}-\eta_{\phi}\bigtriangledown\mathcal{L}_{1}$.
-12:end for
-13$\triangleright$ Generated $p^{p}$ using the updated Encoder.
-14:$p^{p}=SLM_{Encoder}(p)$.
-15$\triangleright$ Generated perturbed rationales from LLM on the server.
-16:$r^{p}=LLM(p^{p})$.
-17$\triangleright$ Decoded perturbed rationales using the updated Encoder-Decoder SLM.
-18:$r=SLM_{Decoder}(r^{p})$.
-19$\triangleright$ Multi-Task Training for Task-Specific SLM based on Datasets $\mathcal{D}$.
-20:for each epoch $t\in[T]$ do
-21:     $\omega^{t+1}\leftarrow\omega^{t}-\eta_{\omega}\bigtriangledown\mathcal{L}_{2}$.
-22:end for
-3.5 Task-Specific SLM Training Module
+3.4 Enhancing Task-Specific SLM via CoT Knowledge Distillation
 In our work, we undertake the training of the client’s Task-Specific SLM tailored for text generation tasks.
@@
 Formally, let’s denote a dataset as $D=\left\{(x_{i},(y_{i},r_{i}))\right\}^{N}_{i=1}$, where $x_{i}$ represents an input, $y_{i}$ represents the associated expected output label, and $r_{i}$ is the corresponding desired rationale.
-We conceptualize learning with rationales as a multi-task learning problem, as illustrated in Figure 5.
+We conceptualize learning with rationales as a multi-task learning problem.
 Specifically, we train the model $f_{\omega}(x_{i})\to(y_{i},r_{i})$ to accomplish not just the prediction of task labels but also the generation of the corresponding rationales based on textual inputs.
@@
 By doing so, we enhance the transparency and explainability of the model.
-The multi-task learning objective can be formulated as follows:
-$\mathcal{L}_{2}=\beta\mathcal{L}_{\text{Label}}+(1-\beta)\mathcal{L}_{\text{Rationale}}$ (6)
+The multi-task learning objective for the Task-Specific SLM can be formulated as follows:
+$\mathcal{L}_{2}=\mathcal{L}_{\text{Label}}+\mathcal{L}_{\text{Rationale}}$ (7)
 where $\mathcal{L}_{\text{Label}}$ is the label prediction loss:
@@
 $\mathcal{L}_{\text{Rationale}}(\omega;\mathcal{D})=\mathbb{E}_{(x,r)\sim\mathcal{D}}\ell_{\text{CE}}(f_{\omega}(x),r)$ (9)
-where $\ell_{\text{CE}}$ is the cross-entropy loss, $f_{\omega}(.)$ is the Task-Specific SLM model, and $\beta$ is the hyperparameters that control the weight of label prediction loss and rationale generation loss.
-Figure 5: Task-Specific SLM Training Module.
+where $\ell_{\text{CE}}$ is the cross-entropy loss, $f_{\omega}(.)$ is the Task-Specific SLM model.
+3.5 Privacy Analysis of FedCoT
+The privacy-protection strategies in FedCoT implement a token-level Exponential Mechanism in feature space, adhering to the $\epsilon$-DP principles.
+This mechanism provides mathematically provable privacy guarantees at the token-level granularity, as extensively validated in privacy-preserving NLP research yue2021differential; chen-etal-2023-customized; tong2025inferdpt.
+Our experimental results further validate this approach: when privacy budget is low, the rationales generated from perturbed prompts show significantly lower similarity to those from original prompts, demonstrating the effectiveness of our privacy protection while acknowledging the inherent privacy-utility trade-off.
 4 Experiments
 4.1 Setup
-We have established a scenario to evaluate the performance of the PDSS framework across a range of text generation tasks.
+We have established a scenario to evaluate the performance of the FedCoT framework across a range of text generation tasks.
 This setup involves a client-server architecture, where the client holds two downstream SLMs: an Encoder-Decoder SLM, which specializes in encoder-decoder functionalities and a Task-Specific SLM, tailored for specific tasks.
 On the server-side, we host a LLM for more general and powerful text generation capabilities.
-Specifically, we have chosen Qwen-14BBai et al.
-(2023) as LLM, while both SLMs are Qwen-0.5BBai et al.
-(2023).
-Table 1 outlines the detailed configurations of both the LLM and the SLMs.
-Setting Server Client Client
-Model Type LLM Encoder-Decoder SLM Task-Specific SLM
-Model Name Qwen-14B Qwen-0.5B Qwen-0.5B
-Parameters(Billion) 14 0.5 0.5
-Table 1: LLM and SLMs Setting of PDSS.
+Specifically, Table 1 outlines the detailed configurations of both the LLM and the SLMs.
+In our experimental setup, the Encoder-Decoder SLM and Task-Specific SLM are the identical architecture.
+SLM
+Setting LLM Encoder-Decoder Task-Specific
+Setting 1 LLaMA3 70B Pythia-1.4B Pythia-1.4B
+Setting 2 Qwen1.5-14B Qwen1.5-0.5B Qwen1.5-0.5B
+Table 1: LLM and SLMs Setting of FedCoT.
 Datasets and Evaluation Metrics.
-We conduct a comprehensive evaluation of PDSS on 4 QA datasets.
-Specifically, we include CommonsenseQA(CQA) Talmor et al.
-(2018), OpenBookQA(OBQA) Mihaylov et al.
-(2018), BoolQ Clark et al.
-(2019), ARC-EClark et al.
-(2018).
+We conduct an evaluation of FedCoT on 4 QA datasets.
+Specifically, we include CommonsenseQA (CQA) talmor2019commonsenseqa, OpenBookQA (OBQA)  mihaylov2018can, BoolQ  clark2019boolq, ArcE clark2018think.
 For these datasets, we primarily use Accuracy as the evaluation metric.
+It’s worth noting that in our experiments, all methods undergo zero-shot evaluation except FewShot(1-shot), and we use the lm-evaluation-harness package eval-harness.
 Baselines.
-Since we incorporate two distinct strategies in the prompt encoder module and rationales decoder module, we denote PDSS method with the Exponential Mechanism Strategy as PDSS-EM and PDSS method with the Encoder-Decoder Strategy as PDSS-ED.
-We conduct a comparative analysis to evaluate the performance of our PDSS framework, which comprises both PDSS-EM and PDSS-ED.
+Since we incorporate two distinct strategies in the prompt encoder and perturbed rationales decoder, we denote FedCoT method with the Exponential Mechanism Strategy as FedCoT-E and FedCoT method with the Adaptive Exponential Mechanism Strategy as FedCoT-A.
+We conduct a comparative analysis to evaluate the performance of our FedCoT framework, which comprises both FedCoT-E and FedCoT-A.
 These baselines included:
@@
   • FewShot-SLM, which represents the few-shot performance of SLM on the client;
-  • Standalone, where the client independently fine-tunes its local model using its own private dataset;
-  • DSSHsieh et al. (2023), where the client fine-tunes its local model by distilling step-by-step LLM method without privacy-preserving.
-4.2 Overall Performance Evaluation
-In this section, we undertake a comprehensive analysis of the task performance of PDSS.
-We assess both the PDSS-EM and PDSS-ED methods against other baselines on Task-Specific SLM across various privacy budgets, denoted by $\epsilon$.
-The results, as presented in Table 2, clearly illustrate that both PDSS-EM and PDSS-ED exhibit significantly better performance when compared to FewShot-SLM and Standalone methods.
-With an increase in the privacy budget $\epsilon$, both the performance of PDSS-EM and PDSS-ED have risen notably.
-Furthermore, PDSS-ED demonstrates notably superior performance compared to PDSS-EM under the same privacy budget $\epsilon$ Specifically, under a privacy budget of $\epsilon=3$, PDSS-EM surpasses the Standalone method by 3.4% and 17% in the CQA and OBQA datasets, respectively, while PDSS-ED outperforms it by 5.2% and 22.4%.
-Similarly, when the privacy budget is increased to $\epsilon=10$, PDSS-EM exceeds the Standalone approach by 6.3% and 21.6% within the CQA and OBQA datasets, respectively, and PDSS-ED beats it by 7.2% and 28.6%.
-Remarkably, across all datasets evaluated, when the privacy budget is set to $\epsilon=10$, PDSS achieves comparable performance to DSS, highlighting its efficacy and versatility in balancing privacy and utility.
-Method CQA OBQA BoolQ ARC-E
+  • Standalone, where the client fine-tunes its local model using its own private dataset;
+  • Non-Private, where the client send its raw local prompt to server, get rationales from LLM and fine-tunes its local model like FedCoT, but without privacy-preserving.
+4.2 Main Results
+In this section, we undertake a comparative analysis of the task performance of FedCoT.
+We assess both the FedCoT-E and FedCoT-A methods against other baselines on Task-Specific SLM under the privacy budget $\epsilon=3$.
+Our experiments encompass two model configurations: Setting 1 (LLM: LLaMA3-70B, Encoder-Decoder SLM & Task-Specific SLM: Pythia-1.4B) and Setting 2 (LLM: Qwen1.5-14B, Encoder-Decoder SLM & Task-Specific SLM: Qwen1.5-0.5B).
+The results, as presented in Table 2, clearly illustrate that both FedCoT-E and FedCoT-A exhibit significantly better performance when compared to FewShot-SLM and Standalone methods.
+Furthermore, FedCoT-A demonstrates notably superior performance compared to FedCoT-E.
+Specifically, take the model Setting 1 as an example, FedCoT-E surpasses the Standalone method by 4.3%, 3.2%, 7.1%, and 5.1% in the CQA, OBQA, BoolQ, and ArcE datasets, respectively.
+Meanwhile, FedCoT-A demonstrates even greater superiority, exceeding the Standalone method by 5.7%, 4.6%, 6.7%, and 6% across the same datasets.
+Model Method CQA OBQA BoolQ ArcE
+Setting 1
+FewShot-LLM 70.29 80.66 90.08 82.69
+FewShot-SLM 21.19 26.60 52.11 28.91
+Standalone 42.43 38.73 73.07 40.33
+Non-Private 49.22 46.07 80.61 48.01
+FedCoT-E 46.70 41.93 80.02 45.42
+FedCoT-A 48.10 43.30 79.77 46.34
+Setting 2
 FewShot-LLM 80.9 82.8 85.2 80.3
@@
 Standalone 55.7 43.4 78.4 50.3
-DSS 59.3 55.1 80.5 57.6
-PDSS-EM($\epsilon=1$) 57.7 49.2 80.1 52.3
-PDSS-EM($\epsilon=3$) 57.6 50.8 79 52.6
-PDSS-EM($\epsilon=5$) 58.8 53.2 80 55.3
-PDSS-EM($\epsilon=10$) 59.2 52.8 80.2 56.2
-PDSS-ED($\epsilon=1$) 58.2 50.8 80.3 56.4
-PDSS-ED($\epsilon=3$) 58.6 53.1 80.2 56.5
-PDSS-ED($\epsilon=5$) 58.3 53.4 80.4 56.3
-PDSS-ED($\epsilon=10$) 59.7 55.8 80.7 57.9
-Table 2: We compare the performance of Task-Specific SLM trained with PDSS-EM and PDSS-ED across different privacy budgets $\epsilon$ against the Task-Specific SLM trained using baseline methods.
-4.3 Reducing Training Data Evaluation
-In this section, we conduct an in-depth analysis to explore the influence of training data size on model performance.
-We compare the PDSS method with the Standalone approach, varying the amount of training data used.
-Table 3 provides a clear illustration of how PDSS(with $\epsilon=3$) consistently outperforms the Standalone method.
-Remarkably, PDSS achieves superior performance even with significantly fewer training samples compared to Standalone.
-More specifically, when trained on merely 75% of the complete CQA, OBQA, and BoolQ datasets, both PDSS-EM and PDSS-ED surpasses the performance of Standalone fine-tuning that has been trained on the entirety of these datasets.
-Likewise, by using only 50% of the full ARC-E dataset, PDSS-EM exceeds the results achieved by Standalone fine-tuning on the complete dataset.
-Furthermore, PDSS-ED exhibits significantly better performance than PDSS-EM across various dataset sizes (ranging from 25% to 100%).
-The results indicate that PDSS is capable of extracting more valuable information from smaller datasets, making it a promising approach in data-scarce environments.
-Task Method 25% 50% 75% 100%
-CQA PDSS-EM 49 53.5 56.7 57.6
-PDSS-ED 54.2 54.6 56.1 58.6
-Standalone - - - 55.7
-OBQA PDSS-EM 34.8 42.2 45.6 50.8
-PDSS-ED 41.4 43.6 50.6 53.1
-Standalone - - - 44.2
-BoolQ PDSS-EM 63 74 78.7 79
-PDSS-ED 72.8 77.6 79.1 80.2
-Standalone - - - 78.4
-ARC-E PDSS-EM 45.3 52.2 53.1 53.8
-PDSS-ED 48 49.7 55.9 56.5
-Standalone - - - 50.3
-Table 3: We compare the performance of Task-Specific SLM trained with PDSS-EM($\epsilon=3$) and PDSS-ED($\epsilon=3$) against Standalone, across a range of dataset sizes from 25% to 100%.
-The ’-’ indicates a method does not apply to the corresponding dataset sizes.
-4.4 Perturbed Rationales Evaluation
-In this section, we focus on analyzing the quality of the perturbed rationales($r^{p}$) generated from the perturbed prompt of LLM based on PDSS-EM and PDSS-ED methods and compare them with the rationales($r$) generated from raw prompt of the LLM.
+Non-Private 59.3 55.1 80.5 57.6
+FedCoT-E 57.6 50.8 79 52.6
+FedCoT-A 58.6 53.1 80.2 56.5
+Table 2: We compare the performance of Task-Specific SLM trained with FedCoT-E ($\epsilon=3$) and FedCoT-A ($\epsilon=3$) against the Task-Specific SLM trained using baseline methods.
+We consider two model settings: Setting 1 (LLM: LLaMA3-70B, Encoder-Decoder SLM & Task-Specific SLM: Pythia-1.4B) and Setting 2 (LLM: Qwen1.5-14B, Encoder-Decoder SLM & Task-Specific SLM: Qwen1.5-0.5B)
+4.3 Performance Evaluation on various SLMs
+In this section, we extend the evaluation of FedCoT’s effectiveness to encompass various client-side SLMs.
+These SLMs include LLaMA2-1.3B xia2024sheared, Qwen1.5-1.8B bai2023qwen, and OPT-1.3B zhang2022opt.
+We have chosen LLaMA3-70B dubey2024llama as LLM.
+Table 3 provides a clear illustration of how FedCoT(with $\epsilon$ = 3) consistently outperforms the Standalone method across various SLMs.
+Dataset Method LLaMA2 Qwen1.5 OPT
+CQA Standalone 61.5 57.8 56.42
+FedCoT-E 63.03 60.30 57.55
+FedCoT-A 64.27 62.21 60.18
+OBQA Standalone 47.53 52.60 40.93
+FedCoT-E 51.73 56.40 49.13
+FedCoT-A 49.8 57.20 48.4
+BoolQ Standalone 81.65 81.41 72.84
+FedCoT-E 83.94 82.59 82.46
+FedCoT-A 82.99 82.90 82.68
+ArcE Standalone 40.33 55.58 45.92
+FedCoT-E 54.11 61.07 49.67
+FedCoT-A 54.66 62.43 50.69
+Table 3: We compare the performance of Task-Specific SLMs, which have been trained with FedCoT-E($\epsilon=3$) and FedCoT-A($\epsilon=3$), against Standalone across various SLMs, including LLaMA2-1.3B, Qwen1.5-1.8B and OPT-1.3B.
+4.4 Ablation Study
+Influence of Privacy Budgets.
+We delve into the influence of privacy budgets on the performance of FedCoT.
+To ensure experimental consistency, we fix the model configuration to Setting 1 (as detailed in Table 1) for all subsequent ablation experiments.
+Table 4 presents an overview of FedCoT’s performance across a range of privacy budgets ($\epsilon=1,3,5,10$).
+As the privacy budget $\epsilon$ increases, the performance of both FedCoT-E and FedCoT-A exhibits a notable uptick.
+Moreover, FedCoT-A consistently outperforms FedCoT-E under identical privacy budget conditions ($\epsilon$).
+When compared alongside Table 2, it becomes evident that with a privacy budget escalated to $\epsilon=10$, FedCoT-E surpasses the Standalone method by 5.6%, 6.1%, 6.3%, and 6.8% within the CQA, OBQA, BoolQ, and ArcE datasets, respectively.
+Similarly, FedCoT-A outperforms it by 4.3%, 7.1%, 6.8%, and 7%.
+Notably, across all evaluated datasets, at a privacy budget of $\epsilon=10$, FedCoT attains performance levels comparable to Non-Private approaches, underscoring its proficiency and adaptability in striking a balance between privacy and utility.
+Method $\epsilon$ CQA OBQA BoolQ ArcE
+FedCoT-E 1 45.63 42.13 78.91 44.84
+3 46.70 41.93 80.02 45.42
+5 46.50 43.35 80.17 46.70
+10 48.03 44.87 79.37 47.14
+FedCoT-A 1 47.31 43.20 79.63 46.65
+3 48.10 43.30 79.77 46.34
+5 47.96 44.20 79.91 48.08
+10 47.74 45.81 79.86 47.30
+Table 4: Comparison of the performance of Task-Specific SLM trained with FedCoT-E and FedCoT-A across different privacy budgets $\epsilon$.
+Influence of Perturbed Rationales Decoding.
+We undertake an analysis to investigate the effects of perturbed rationales decoding on FedCoT when $\epsilon=3$.
+Table 5 offers a comparison of FedCoT’s performance, contrasting the results when perturbed rationales decoding is employed (FedCoT-E w/ and FedCoT-A w/) versus when it is not (FedCoT-E w/o and FedCoT-A w/o).
+Specifically, FedCoT-E w/ surpasses the FedCoT-E w/o by 2%, 1.3%, 1.5%, and 0.6% in the CQA, OBQA, BoolQ, and ArcE datasets, respectively.
+Meanwhile, FedCoT-A w/ demonstrates even greater superiority, exceeding the FedCoT-A w/o by 1.8%, 1.6%, 0.7%, and 3% across the same datasets.
+The findings unequivocally demonstrate that FedCoT exhibits superior performance when perturbed rationales decoding is utilized, as compared to when it is absent.
+Decoding
+Method Dataset w/ w/o
+FedCoT-E CQA 46.70 44.79
+OBQA 41.93 40.6
+BoolQ 80.02 78.5
+ArcE 45.42 44.78
+FedCoT-A CQA 48.10 46.26
+OBQA 43.30 41.7
+BoolQ 79.77 79.06
+ArcE 48.08 45.13
+Table 5: Comparison of Task-Specific SLM Performance in FedCoT: With vs. Without perturbed rationales decoding.
+Perturbed Rationales vs Original Rationales.
+We focus on analyzing the quality of the perturbed rationales ($r^{p}$) generated from the perturbed prompt of LLM based on FedCoT-E and FedCoT-A methods and compare them with the rationales ($r$) generated from raw prompt of the LLM.
 To evaluate the similarity between $r^{p}$ and $r$, we use TokenRatio metric.
 A higher TokenRatio indicates a greater degree of similarity between the perturbed and original rationales.
-For more details about TokenRatio, please refer to Appendix C.
-As shown in Table 4, with an increase in the privacy budget $\epsilon$ and a corresponding decrease in perturbation, both the TokenRatio of PDSS-EM and PDSS-ED have risen notably.
-Furthermore, in most of tasks, the TokenRatio of PDSS-ED is higher than that of PDSS-EM in the same level of privacy budget $\epsilon$.
-The experimental results confirm that the TokenRatio observed in the perturbed rationales produced by both PDSS-EM and PDSS-ED, positively correlate with the privacy budget $\epsilon$.
+TokenRatio($r^{{}^{\prime}},r$).
+This metric calculates the unique words($u$) in $r^{{}^{\prime}}$ and counts how many of these words are also present in $r$, denoted as $i$.
+The TokenRatio is then calculated as $i$ divided by the total number of unique words in $r^{{}^{\prime}}$ ($|u|$).
+As shown in Table 6, with an increase in the privacy budget $\epsilon$ and a corresponding decrease in perturbation, both the TokenRatio of FedCoT-E and FedCoT-A have risen notably.
+Furthermore, in most of tasks, the TokenRatio of FedCoT-A is higher than that of FedCoT-E in the same level of privacy budget $\epsilon$.
+The experimental results confirm that the TokenRatio observed in the perturbed rationales produced by both FedCoT-E and FedCoT-A, positively correlate with the privacy budget $\epsilon$.
 This suggests that as the privacy constraints are relaxed (higher $\epsilon$ values), the perturbed rationales become more similar to the original rationales.
-This finding is significant as it demonstrates the trade-off between privacy protection and the utility of the generated rationales.
-Method CQA OBQA BoolQ ARC-E
-PDSS-EM($\epsilon=1$) 19.8 26.2 26.6 24.6
-PDSS-EM($\epsilon=3$) 29.2 37.2 35.5 33.9
-PDSS-EM($\epsilon=5$) 48.8 59.6 55.2 53.9
-PDSS-EM($\epsilon=10$) 69.7 72 74.6 68.2
-PDSS-ED($\epsilon=1$) 26.7 33.1 29.7 31
-PDSS-ED($\epsilon=3$) 33.1 40.9 40.4 42.9
-PDSS-ED($\epsilon=5$) 49.6 61 57.5 63.5
-PDSS-ED($\epsilon=10$) 57.2 68.3 68 74.2
-Table 4: We conduct a comparative analysis to assess the perturbed rationales produced by PDSS-EM and PDSS-ED methods against the original, unperturbed (raw) rationales that are directly generated from the raw prompt of the LLM.
-4.5 Decoded Rationales Evaluation
-In this section, we delve into the quality analysis of the decoded rationales produced by the rationales decoder module based on PDSS-EM and PDSS-ED methods.
+Method $\epsilon$ CQA OBQA BoolQ ArcE
+FedCoT-E 1 23.8 33 34.5 26.7
+3 30.8 45.26 48.5 44.7
+5 43.2 66.3 72.8 67.4
+10 48.5 75.8 85.4 74.5
+FedCoT-A 1 34.5 37.9 47.1 20.7
+3 34.5 49.5 59.6 30
+5 45.2 69.6 77.4 36.2
+10 48.6 76.12 84.2 38.6
+Table 6: We conduct a comparative analysis to assess the perturbed rationales produced by FedCoT-E and FedCoT-A methods against the original rationales that are directly generated from the raw prompt of the LLM.
+Metric used: TokenRatio.
+Decoded Rationales vs Original Rationales.
+We delve into the quality analysis of the decoded rationales produced by the rationales decoder module based on FedCoT-E and FedCoT-A methods.
 We compare these decoded rationales against those generated directly from raw prompt of the LLM.
 We utilize the TokenRatio metric to assess their similarities.
-As shown in Table 5, in contrast to FewShot-SLM, it becomes apparent that the decoded rationales’ quality based on PDSS-EM and PDSS-ED methods isn’t solely reliant on the locally decoded SLM.
+As shown in Table 7, in contrast to FewShot-SLM, it becomes apparent that the decoded rationales’ quality based on FedCoT-E and FedCoT-A methods isn’t solely reliant on the locally decoded SLM.
 The perturbed rationales crafted by the LLM indeed fulfill their intended purpose.
-When juxtaposed with Table 4, it’s clear that at comparable $\epsilon$ levels, the TokenRatio for the decoded rationales surpass those of the perturbed rationales in the PDSS-EM and PDSS-ED methods.
-This underscores the effectiveness of the rationales decoder module in the PDSS-EM and PDSS-ED methods.
-Furthermore, with the increase of the privacy budget $\epsilon$, the TokenRatio for the decoded rationales generated by both PDSS-EM and PDSS-ED have increased significantly.
-This suggests that as the privacy constraints are relaxed (higher $\epsilon$ values), the decoded rationales become more similar to the original rationales.
-For more details about comparative analysis of perturbed rationales and decoded rationales, please refer to Appendix D.
-Method CQA OBQA BoolQ ARC-E
-FewShot-SLM 43.3 43.4 51.9 42.6
-PDSS-EM($\epsilon=1$) 38.3 37.1 38.4 41.5
-PDSS-EM($\epsilon=3$) 41.9 41.3 41.7 45.6
-PDSS-EM($\epsilon=5$) 53.1 54 55 58.3
-PDSS-EM($\epsilon=10$) 71.1 63 73.6 70.4
-PDSS-ED($\epsilon=1$) 57.2 53.4 45.2 57.5
-PDSS-ED($\epsilon=3$) 59 55.1 48 59.4
-PDSS-ED($\epsilon=5$) 59.8 59.5 55.7 65.5
-PDSS-ED($\epsilon=10$) 62 62.3 63.4 70.1
-Table 5: We conduct a comparative analysis to assess the decoded rationales produced by PDSS-EM and PDSS-ED methods against the original, unperturbed (raw) rationales that are directly generated from the raw prompt of the LLM.
+When juxtaposed with Table 6, it’s clear that at comparable $\epsilon$ levels, the TokenRatio for the decoded rationales surpass those of the perturbed rationales in the FedCoT-E and FedCoT-A methods.
+This underscores the effectiveness of the rationales decoder module in the FedCoT-E and FedCoT-A methods.
+Method $\epsilon$ CQA OBQA BoolQ ArcE
+FewShot-SLM - 42.9 54.5 35.8 28.6
+FedCoT-E 1 36 46.33 44.13 32.7
+3 39 53.77 53.1 46
+5 44.8 67.9 73.9 60.1
+10 48.4 75.1 85.4 66.7
+FedCoT-A 1 41.1 60.36 62.8 42.19
+3 45.8 65.35 64.7 42.99
+5 50 75.5 72.9 44.3
+10 53.3 78.9 76.6 45.3
+Table 7: We conduct a comparative analysis to assess the decoded rationales produced by FedCoT-E and FedCoT-A methods against the original rationales that are directly generated from the raw prompt of the LLM.
+Metric used: TokenRatio.
+Outperforming Standalone with 50% Data.
+We conduct an in-depth analysis to explore the influence of training data size on model performance.
+We compare the FedCoT method with the Standalone approach, varying the amount of training data used.
+Table 8 provides a clear illustration of how FedCoT(with $\epsilon=3$) achieves superior performance even with significantly fewer training samples compared to Standalone.
+More specifically, when trained on merely 50% of the complete CQA, OBQA, BoolQ, and ArcE datasets, both FedCoT-E and FedCoT-A either surpass or closely match the performance of Standalone method.
+Dataset Method 25% 50% 75% 100%
+CQA FedCoT-E 37.74 42.63 44.56 46.7
+FedCoT-A 39.28 44.77 44.00 48.1
+Standalone - - - 42.43
+OBQA FedCoT-E 32.4 38.27 40.67 41.93
+FedCoT-A 34.07 38.08 42.00 43.3
+Standalone - - - 38.73
+BoolQ FedCoT-E 69.96 72.26 77.67 80.02
+FedCoT-A 69.61 73.73 77.82 79.77
+Standalone - - - 73.07
+ArcE FedCoT-E 37.79 41.42 42.22 45.42
+FedCoT-A 37.64 41.86 45.28 46.34
+Standalone - - - 40.33
+Table 8: We compare the performance of Task-Specific SLM trained with FedCoT-E($\epsilon=3$) and FedCoT-A($\epsilon=3$) against Standalone, across a range of dataset sizes from 25% to 100%.
+The ’-’ indicates a method does not apply to the corresponding dataset sizes.
 5 Conclusions
-We introduced PDSS, a privacy-preserving framework for LLM distillation, addressing domain-specific knowledge privacy and resource constraints.
-PDSS employs a server-client architecture with prompt encoding, rationale generating, rationale decoding, and task-specific SLM training, bridging the gap between LLM and SLM while maintaining data privacy.
-Experiments on various text generation tasks demonstrate PDSS’s ability to enhance SLM performance with LLM support while prioritizing data privacy.
+In this study, we introduce FedCoT, a federated framework designed to distill knowledge from LLMs to SLMs in resource-constrained environments.
+FedCoT facilitates secure knowledge transfer from LLMs to SLMs by leveraging perturbed prompts and rationales, thereby enhancing the performance of SLMs without compromising user privacy.
+We present two innovative privacy protection strategies, including an Adaptive Exponential Mechanism strategy, which effectively balance privacy preservation and the usability of rationales.
+Experiments on various text generation tasks demonstrate FedCoT’s ability to enhance SLM performance with LLM support while prioritizing data privacy.
 Limitations
-Our current study faces limitations due to computational and storage constraints, which hinder our ability to experiment with larger model sizes.
-Additionally, our evaluation of PDSS has been restricted to the Qwen model architecture, leaving the possibility that PDSS may need to be further explored in other model architectures.
-We intend to tackle these issues in future research endeavors.
-Appendix A Rationales Generation through COT
-We utilize the rationales data generated by server-side LLM through chain-of-thought (CoT)Wei et al.
-(2022)Hsieh et al.
-(2023) technique to enhance the performance of the client’s task-specific SLM.
+While our proposed FeCoT framework demonstrates promising results in privacy-preserving knowledge transfer from LLMs to SLMs, it is important to acknowledge several considerations that could be addressed in future work.
+Firstly, the framework’s performance benefits are contingent upon the server-side LLM’s CoT reasoning capabilities.
+Although contemporary LLMs like GPT-4 and LLaMA exhibit strong reasoning skills, frameworks such as FedCoT may encounter limitations when deployed with less sophisticated LLMs.
+This suggests an opportunity for further research to enhance FedCoT’s robustness against variability in LLM reasoning abilities.
+Secondly, our evaluation primarily focused on LLaMA and Qwen as the server-side LLMs, with client-side SLMs including Pythia, LLaMA, Qwen, and OPT.
+While these models are representative of current state-of-the-art architectures, extending testing to a more diverse set of LLMs could provide deeper insights into FedCoT’s generalizability.
+Appendix A FedCoT’s Computational and Communication Overhead
+FedCoT is designed to be efficient and scalable in resource-constrained environments.
+The communication overhead is minimal, with costs comparable to plaintext data transmission.
+Computational requirements are equivalent to standard SLM fine-tuning (SFT) on local tasks.
+Our experimental validation was conducted using NVIDIA V100 GPUs, demonstrating practical deployment feasibility.
+Appendix B Rationales Generation through CoT
+We utilize the rationales data generated by server-side LLM through chain-of-thought (CoT)wei2022chainhsieh2023distilling technique to enhance the performance of the client’s task-specific SLM.
 These rationales justify the predicted labels and serve as insightful guidance for training smaller and domain-specific models.
@@
 Therefore, we harness these rationales as enriched training material for small language models, employing a multi-task training paradigm that encompasses both label prediction task and rationale prediction task.
-Appendix B More on Experimental Details
-B.1 Hyperparameter Settings
+Appendix C More on Experimental Details
+C.1 Hyperparameter Settings
 SLM Parameters.
@@
 Furthermore, the learning rates for $\eta_{\phi}$ and $\eta_{\omega}$ were established at 5e-5.
-B.2 Data Splitting
-For the datasets CQA/OBQA/BoolQ//ARC-E/, all splits (training, validation, and test) were downloaded from HuggingFace Lhoest et al.
-(2021).
+C.2 Data Splitting
+For the datasets CQA/OBQA/BoolQ//ArcE/, all splits (training, validation, and test) were downloaded from HuggingFace lhoest-etal-2021-datasets.
 During the training of the Encoder-Decoder SLM, we randomly divided the training data into two equal parts.
 One part was designated as the public dataset, while the other part was allocated as the private dataset for the client.
-B.3 Dataset Licenses
-For the datasets CQA/OBQA/BoolQ//ARC-E/ were downloaded from HuggingFaceLhoest et al. (2021) and under Apache License, Version 2.0.
-B.4 Machine Configuration
-The experiments were conducted on machines equipped with 4 Nvidia V100 32G.
-Appendix C The Definition of TokenRatio Metric
-TokenRatio($r^{{}^{\prime}},r$).
-This metric calculates the unique words($u$) in $r^{{}^{\prime}}$ and counts how many of these words are also present in $r$, denoted as $i$.
-The TokenRatio is then calculated as $i$ divided by the total number of unique words in $r^{{}^{\prime}}$ ($|u|$).
-Figure 6: Comparative Analysis of Perturbed Rationales and Decoded Rationales.
-Appendix D Comparative Analysis of Perturbed Rationales and Decoded Rationales
-As shown in Figure 6, we conduct a comparison of the quality between the perturbed rationales and the decoded rationales, employing both the PDSS-EM and PDSS-ED methods across various privacy budgets denoted by $\epsilon$.
-For clarity, we designate the perturbed rationales generated using the PDSS-EM and PDSS-ED methods as P-PDSS-EM and P-PDSS-ED, respectively.
-Similarly, the decoded rationales derived from these methods are denoted as D-PDSS-EM and D-PDSS-ED.
-It’s clear that at comparable $\epsilon$ levels, the TokenRatio for decoded rationales consistently surpasses that of perturbed rationales in most tasks, when utilizing the PDSS-EM and PDSS-ED methods.
-This finding underscores the remarkable effectiveness of the rationales decoder module within both the PDSS-EM and PDSS-ED frameworks.
+C.3 Dataset Licenses
+For the datasets CQA/OBQA/BoolQ//ARC-E/ were downloaded from HuggingFacelhoest-etal-2021-datasets and under Apache License, Version 2.0.
+C.4 Machine Configuration
+The experiments were conducted on machines equipped with 4 and 8 NVIDIA V100 32G.